One cleaning fluid contains 8% ammonia, while a second contains 18% ammonia. In mixing these fluids, how many liters of the second needs to be added to 15 liters of the first, for the percent of the ammonia in the mixture to be 14%?
Answer by lwsshak3(11628)

One cleaning fluid contains 8% ammonia, while a second contains 18% ammonia. In mixing these fluids, how many liters of the second needs to be added to 15 liters of the first, for the percent of the ammonia in the mixture to be 14%?
let x= liters of the 18% (second fluid) need to be added to the 15 liters of the 8% fluid
x+15 = liters of the 14% mixture
the sum of ammonia in the two fluids=amount of ammonia in the mixture
.18x+.08*15=.14(x+15)
multiply by 100
18x+8*15=14(x+15)
18x+120=14x+210
4x=90
x=90/4=22.5 liters
ans: 22.5 liters of the second fluid must be added to 15 liters of the first fluid to get a mixture of 14% ammonia